OpenStudy (dls):

Let f(x)=|x| and g(x)=[x]. Evaulate:

OpenStudy (dls):

(fog)(-5/3)-(gof)(-5/3)

OpenStudy (dls):

isn't @ParthKohli enough for this? :|

OpenStudy (dls):

I am getting zero :(

OpenStudy (dls):

@terenzreignz @amistre64

OpenStudy (anonymous):

it must be one :)

OpenStudy (dls):

how?

Parth (parthkohli):

What is [x]? Like \(\lfloor x \rfloor\)

OpenStudy (dls):

gINT

OpenStudy (anonymous):

fog(x) = |[x]| fog(-5/3) = fog(-1.67) the greatest integer function for -1.67 gives -2.. taking absolute value.. we get fog(-5/3) = 2 gof(x) = [|x|] gof(-5/3) = fog(-1.67 the absolute value will be 1.67 and it's greatest integer function value will be 1.. So 2-1 = 1 :D

OpenStudy (dls):

[1.67] should be 1..

OpenStudy (dls):

1-1 is 0

OpenStudy (anonymous):

but its negative of 1.67 man !

OpenStudy (dls):

oh yeah!sorryy!

OpenStudy (dls):

okay,got it :P